OpenStudy (anonymous):

18. Use the given degree of confidence and sample data to construct a confidence interval for the population proportion p. n = 110, x = 55; 88% confidence

OpenStudy (anonymous):

First solve for p = x/n = 55/110 = 0.5 then find the error, \[E = (z _{a/2})\sqrt{(p(1-p))/n} = (1.555)\sqrt{((0.5)(0.5))/110} = 0.078\] The 88% CIE is (0.5-0.078 , 0.5+0.078) = (0.422 , 0.578)